Can someone help me on this problem?
I need to perform the indicated operations on:
6 2/9 + 1 3/4 - 3 5/6
Thanks,
Lyn

Found 2 solutions by fractalier, Earlsdon:
Answer by fractalier(6550)

Combining fractions with unlike denominators requires changing them into common, like denominators first...the lowest common denominator here is 36...so we have
6 2/9 + 1 3/4 - 3 5/6 =
6 8/36 + 1 27/36 - 3 30/36 =
7 35/36 - 3 30/36 =
4 5/36

Answer by Earlsdon(6294)

Here's one way to approach this problem.
1) You can convert the given mixed numbers to improper fractions, find their common denominator, then perform the indicated operations.
6 2/9 = 56/9
1 3/4 = 7/4
3 5/6 = 23/6
Now you need to find the common denominator of 9, 4, and 6.
What number will 9, 4, and 6 go into?
Answer: 36
Now convert the improper fractions to 36ths.
56/9 = 224/36 (You multiplied top and bottom by 4)
7/4 = 63/36 (You multiplied top and bottom by 9)
23/6 = 138/36 (You multiplied top and bottom by 6)
Now perform the indicated operations.
224/36 + 63/36 - 138/36 = 287/36 - 138/36 = 149/36 Now simplify this.
149/36 = 4 5/36